# Infinite-component $BF$ topological field theory: Nexus of fracton order, Toeplitz braiding, and non-Hermitian Amplification

### v2

@@ V Conclusion and outlook
Several directions merit future exploration. A natural next step is to construct explicit lattice realizations of iBFBF theories that support Toeplitz braiding, enabling direct numerical studies and potentially guiding experimental implementations. Conversely, it is equally worthwhile to develop field-theoretical approaches, including iBFBF theory, to revisit topics extensively studied in lattice settings—such as generalized entanglement renormalization [16] and foliation structures [5, 102, 103]—in the context of four-dimensional fracton topological orders. In this work we focused on particle–loop braiding arising from stacking pure $BF$ theories. Extending the construction by incorporating twisted terms to realize multi-loop and Borromean-rings braiding remains an open problem [142, 134, 130, 135, 153, 162, 163, 164, 190, 191]. Another intriguing direction is to incorporate global or subsystem symmetries into iBFBF theories and to explore the resulting symmetry-fractionalization patterns.
@@ A.2 Detailed calculation on ZSMs of nSSH-type $K$ matrices
Denote the corresponding candidate boundary zero mode as $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$. If $|\beta_{1}|<1$, $|m_{1}|>|n_{1}|$, the following
@@
If $|m_{1}|>|n_{1}|$, the following $\mathbf{u}_{1}$ is a legitimate solution to LZSM. If $|m_{2}|>|n_{2}|$, the following $\mathbf{u}_{2}$ is automatically a legitimate solution to LZSM.
@@
Denote the corresponding candidate boundary zero mode as $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$. If $|\beta_{1}|>1$, $|m_{1}|>|n_{1}|$, the following $\mathbf{v}_{1}$

### v3

@@ V Conclusion and outlook
Several directions merit future exploration. A natural next step is to construct explicit lattice realizations of iBFBF theories that support Toeplitz braiding, enabling direct numerical studies and potentially guiding experimental implementations. Conversely, it is equally worthwhile to develop field-theoretical approaches, including iBFBF theory, to revisit topics extensively studied in lattice settings—such as generalized entanglement renormalization [16] and foliation structures [5, 17, 102]—in the context of four-dimensional fracton topological orders. In this work we focused on particle–loop braiding arising from stacking pure $BF$ theories. Extending the construction by incorporating twisted terms to realize multiloop and Borromean-rings braiding remains an open problem [141, 133, 129, 134, 152, 160, 161, 162, 189, 190]. Another intriguing direction is to incorporate global or subsystem symmetries into iBFBF theories and to explore the resulting symmetry-fractionalization patterns.
@@ A.2 Detailed calculation on ZSMs of nSSH-type $K$ matrices
Denote the corresponding candidate boundary zero mode as $\mathbf{u}_{1}$ and $\mathbf{u}_{2}$. If $|\beta_{1}|<1$, $|m_{1}|>|n_{1}|$, then
@@
If $|m_{1}|>|n_{1}|$, then $\mathbf{u}_{1}$ is a legitimate solution to LZSM. If $|m_{2}|>|n_{2}|$, then $\mathbf{u}_{2}$ is automatically a legitimate solution to LZSM.
@@
Denote the corresponding candidate boundary zero mode as $\mathbf{v}_{1}$ and $\mathbf{v}_{2}$. If $|\beta_{1}|>1$, $|m_{1}|>|n_{1}|$, then $\mathbf{v}_{1}$